Can you please help me solve this problem:
1.) A teacher is 15 years older than her student. In 15 years, the ratio of their ages will be 3:2. Find their present ages.


Answer by ptaylor(2198)

Let x=student's age; x+15=students age in 15 years
Then x+15=teacher's age; x+30=teacher's age in 15 years
Now we are told that:
(x+30)/(x+15)=3/2 multiply each term by 2(x+15) or cross-multiply
2(x+30)=3(x+15) get rid of parens (distributive law)
2x+60=3x+45 subtract 60 and also 3x from each side
2x-3x+60-60=3x-3x+45-60 collect like terms
-x=-15 multiply each side by -1
x=15---------------------------student's age
x+15=15+15=30------------------------teacher's age
CK
(30+15)/(15+15)=3/2
45/30=3/2
3/2=3/2
Hope this helps---ptaylor